If the Ka of a monoprotic weak acid is 2.9×10−6, what is the pH of a 0.31M solution of this acid?
1
Expert's answer
2020-04-19T07:10:13-0400

pH=-lg[H+];

[H+]=√(Ka*C(acid))=√0.31*2.9*10^-6=9.48*10^-4 ;

pH=-lg(9.48*10^-4)=3

Answer: pH=3